OpenStudy (anonymous):

Will medal and fan :) A football quarterback has 2 more chances to throw a touchdown before his team is forced to punt the ball. He misses the receiver on the first throw 30% of the time. When his first throw is incomplete, he misses the receiver on the second throw 10% of the time. What is the probability of not throwing the ball to a receiver on either throw?

OpenStudy (anonymous):

I got the answer 3%. Just looking for confirmation.

OpenStudy (anonymous):

Oh that's fine, let me explain Let A be the event he misses on the first throw. Then P(A) = 0.3. If he misses on the first throw, let B be the event he misses on the second throw. Then P(B|A) = 0.1. P(A∩B)=P(A)×P(B|A)=0.3×0.1

OpenStudy (anonymous):

Thanks for the quick reply! I did the same process too yay! Thanks for your time!